# Variational Control for Guidance in Diffusion Models

### v1

@@ C.1 Evaluation on Distortion Metrics
In this work, we primarily optimize all competing methods for perceptual quality. However, for completeness, we compare the performance of our proposed method with other baselines on recovery metrics like PSNR and SSIM. Tables 11 and 12 compare our proposed method, NDTM, with competing baselines for linear and non-linear inverse problems. We find that NDTM performs on par with other methods for the super-resolution task. However, for random inpainting and non-linear deblur, NDTM outperforms competing methods in terms of distortion metrics like PSNR. Since NDTM also outperforms existing baselines in terms of perceptual quality (see Table 3), our method provides a better distortion-perception tradeoff.
@@ C.2 Runtime
Figure 5: Qualitative comparison between NDTM and competing baselines on the Non-Linear Deblurring task. NDTM better recovers the structure of the image compared to other baselines. We find that DPS is unstable for this task often generating reconstructions with no information.
Figure 6: Qualitative comparison between NDTM and competing baseline (DMPLug) on the blind image deblurring task. NDTM better recovers the details and structure of the image compared to the baseline. We find DMPlug introduces noisy artifacts and blurry images in some samples.
Figure 7: Qualitative comparison between NDTM and competing baselines on the Random Inpainting (90%) Task. NDTM better recovers the structure of the image compared to other baselines.
Figure 8: Qualitative comparison between NDTM and competing baselines on 4x super-resolution task. NDTM better recovers the structure of the image compared to other baselines. Best viewed when zoomed in.

### v2

@@ C.1 Evaluation on Distortion Metrics
In this work, we primarily optimize all competing methods for perceptual quality. However, for completeness, we compare the performance of our proposed method with other baselines on recovery metrics like PSNR and SSIM. Tables 12 and 13 compare our proposed method, NDTM, with competing baselines for linear and non-linear inverse problems. We find that NDTM performs on par with other methods for the super-resolution task. However, for random inpainting and non-linear deblur, NDTM outperforms competing methods in terms of distortion metrics like PSNR. Since NDTM also outperforms existing baselines in terms of perceptual quality (see Table 2), our method provides a better distortion-perception tradeoff.
@@ C.2 Runtime
Figure 6: Qualitative comparison between NDTM and top competing baselines (see Table 1) on the Non-Linear Deblurring task for ImageNet. NDTM better recovers the structure of the image compared to other baselines. Best viewed when zoomed in.
Figure 7: Qualitative comparison between NDTM and top competing baselines (see Table 1) on the Non-Linear Deblurring task for FFHQ. NDTM better recovers the structure of the image compared to other baselines. Best viewed when zoomed in.
Figure 8: Qualitative comparison between NDTM and competing baseline (DMPlug) on the blind image deblurring task (see Table 3). NDTM better recovers the details and structure of the image compared to the baseline. We find DMPlug introduces noisy artifacts and blurry images in some samples. Best viewed when zoomed in.
Figure 9: Qualitative comparison between NDTM and top competing baselines (See Table 2) on the Random Inpainting (90%) Task for ImageNet. NDTM performs better or/par compared to other baselines. Best viewed when zoomed in.
Figure 10: Qualitative comparison between NDTM and top competing baselines (See Table 2) on the Random Inpainting (90%) Task for FFHQ. NDTM performs better or/par compared to other baselines. Best viewed when zoomed in.
Figure 11: Qualitative comparison between NDTM and top competing baselines (See Table 2) on 4x super-resolution task for ImageNet. NDTM performs better or/par compared to other baselines. Best viewed when zoomed in.
Figure 12: Qualitative comparison between NDTM and top competing baselines (See Table 2) on 4x super-resolution task for FFHQ. NDTM performs better or/par compared to other baselines. Best viewed when zoomed in.